### 17 Dec 2023 v3

@@
Summary We investigate long-term cognitive effects of an intervention, where systolic blood pressure (sBP) is monitored at more optimal levels, in a large representative sample. A limitation with previous research on the potential risk reduction of such interventions is that they do not properly account for the reduction of mortality rates. Hence, one can only speculate whether the effect is a result from changes in cognition or changes in mortality. As such, we extend previous research by providing both an etiological and a prognostic effect estimate. To do this we propose a Bayesian semi-parametric estimation approach for an incremental intervention, using the extended G-formula. We also introduce a novel sparsity-inducing Dirichlet hyperprior for longitudinal data, demonstrate the usefulness of our approach in simulations, and compare the performance relative to other Bayesian decision tree ensemble approaches. In our study, there were no significant prognostic- or etiological effects across all ages, indicating that sBP interventions likely do not have a strong effect on memory neither at the population level nor at the individual level.
Key words: BART, dropout, G-computation, L-DART, missing not at random, deaths
@@ 1. Introduction
Two approaches for hypothetical interventions that only intervene on subjects that do not reaches a prespecified threshold are the threshold intervention [35] and the representative intervention (e.g. [27, 41]). Although both of these ensure treatment remains within a prespecified interval, it is common that a significant proportion of the population cannot control their systolic blood pressure (sBP) at optimal levels, due to, for example, poor, or absent, response to treatment (requiring the use of multiple medications). One alternative approach in these settings is a incremental intervention where a subjects treatment instead is shifted downward (or upwards) by some prespecified function [11, 25]. However, their approaches are not valid for longitudinal data with dropout and deaths.
Attrition due dropout and deaths is common in longitudinal aging studies, especially if subjects are followed over a long time period. The G-formula can address both missing response and missing covariate data under an assumption of missing at random (MAR; [32, 41, 18]). Strategies for sensitivity analyses have further been proposed in situations where the missingness is thought to be missing not at random (MNAR) [14, 15]. However, cognition after death is usually not defined and not of interest to study, prompting consideration of a mortal cohort approach [7]. Mortal cohort inferences truncate participants’ outcomes after death. The first approach, partly conditional inference (PCI), condition on the sub-population who are still alive at that time point [19, 38, 37], the second approach, principal stratification, condition on the subpopulation surviving irrespective of treatment [8, 9]. Although the two approaches have similarities, they differ with respect to what type of information they generate. PCI has been proposed for estimating non-causal associations and is particularly useful for prognostication, i.e. when interest is in the health outcome among all individuals who are alive. The approach has for example been considered when intervening on risk factors for coronary heart disease [35]. In contrast, principal stratification is useful in situations when survival differ between treated and controls, and has been applied when interest is in etiological effect estimates [12, 36, 34, 14]. Since the main focus of previous intervention studies was to make prognostic rather than etiological intervention estimates, they have only considered PCI (or simply immortal cohorts) and principal stratification has not been adapted to this setting.
@@
In the present study we investigate long-term cognitive effects of an sBP intervention, using prospective cohort data from a large representative sample [26]. An important limitation in prior research is the lack of sufficient consideration of altered mortality rates in evaluating the potential risk reduction. As such, we propose both partly conditional inference and inference on the principal strata. Thus, our study extends previous research by providing both etiological and prognostic intervention effect estimates, which are both critical for public health practice. In particular, we will be able to answer the following research questions; Would the proposed sBP intervention affect cognition levels in the general population? and Would the proposed intervention affect cognition at the individual level? These will offer a more comprehensive understanding of the intervention’s impact. For estimation, we propose a Bayesian semi-parametric approach based on the extended G-formula [31]. In addition, we introduce a novel sparsity-inducing Dirichlet hyperprior designed for regression with longitudinal data, and demonstrate the usefulness of our approach in a simulation relative to other Bayesian decision tree ensemble approaches.
@@
The paper is organized as follows. In Section 2, we present the Betula data. In Section 3, we introduce the proposed intervention in a Bayesian setting. In Section 4, we describe the two types of mortal cohort inferences. In Section 5, we provide computational details and introduce a sparsity-inducing Dirichlet hyperprior for longitudinal data. In Section 6, we conduct a simulation study to assess the operating characteristics of the new prior, and in Section 7, we apply the methodology to the Betula data to answer the substantive question of interest. We close in Section 8 with some conclusions.
@@ 2 A blood pressure intervention
The subjects in this study are all randomly recruited participants in the Betula prospective cohort study, a population-based study with the objective to study memory, aging and dementia [26]. We consider 15-years of longitudinal data and four waves of data collection, administered in approximately five years intervals, where the first wave was initiated in 1988-1990. A total of n=800 participants were included, 100 participants from each of the eight age-cohorts: $35,40,\ldots,70$ years old at enrollment.
The cognitive outcome in this study was assessed using a composite episodic memory task, with a range of 0 to 76 ([16]). We consider a set of baseline confounding factors; age at enrollment, sex, and years of education, a set of time-varying cardiovascular risk factors; Body mass index (BMI), smoking, diastolic blood pressure, serum-cholesterol, and self reported type 2 diabetes, as well as, previous episodic memory measures. See Table A3 for baseline summary statistics. 38% dropped out of the study before the 15 year follow-up and of these 18% died or developed dementia.
@@
3 Methods
Consider a prospective cohort study where measurements on a set of variables are collected on each of $i=1,2,\ldots,n$ subjects over a specified follow-up period at $t=1,2,\ldots,T$ fixed time points. Let $L_{it}$ denote the vector of possibly time-varying confounders and $A_{it}$ denote a continuous treatment measured at time $t$. Denote the continuous outcome measure by $Y_{it}$, where interest in this study is the outcome measure at the last test wave $T$, i.e. $Y_{iT}$. Furthermore, let $S_{it}=1$ indicate if a subject is alive at time point $t$ and $S_{it}=0$ otherwise. Let $R_{it}=1$ indicate if subject $i$ has completed the testing at time $t$, otherwise $R_{it}=0$. We have monotone missingness, so if $R_{it}=0$, $R_{ik}=0$ for $k>t$. Furthermore, let $T^{r}_{i}$ denote the number of time points a subject participates in the study, and $T^{s}_{i}\geq T^{r}_{i}$ denotes the number of time points the subject is alive. Let the observed history of a time-varying variable be denoted with an overbar and the future with an underbar. We also write $H_{i}$ to denote all the observed past history just prior to the outcome variable at the last time point $T$, such that $H_{i}=(\bar{Y}_{it-1}=\bar{y}_{it-1},\bar{A}_{it}=\bar{a}_{it},\bar{L}_{it}=\bar{l}_{it},\bar{R}_{it}=\bar{r}_{it},\bar{S}_{it}=1)$, with support on the set of possible histories $\mathcal{H}_{i}$ and where lower-case letters denote possible realizations of a random variable. At each time point $t$, we assume data is observed in the temporal ordering $S_{it},R_{it},L_{it},A_{it},Y_{it}$, and for each subject we observe $(\bar{s}_{iT^{s}_{i}},\bar{r}_{iT^{r}_{i}},\bar{l}_{iT^{r}_{i}},\bar{a}_{iT^{r}_{i}},\bar{y}_{iT^{r}_{i}})$. Below, we suppress the subscript $i$ to simplify notation.
3.1 A hypothetical incremental intervention on the NVT
[…]
Here we propose an incremental intervention [11, 25] based on the NVT, where treatment is shifted at time $t$ based on a parameter $\delta_{t}$. In a Bayesian setting it is possible to quantify the uncertainty in the intervention by imposing a prior on the shift parameter $\delta_{t}$, $p^{\delta_{t}}(\cdot)$, reflecting the researchers beliefs about the possible measures for treatment to take under the hypothetical intervention. We describe possible priors in relation to the Betula data in Section 7.
Similar to a threshold intervention [35], we one only intervene on subjects whose NVT at time $t$ does not meet a particular threshold, here denoted by $\tau$. However, instead of assigning treatment to the exact threshold value, $\tau$, treatment is assigned using the prior, $p^{\delta_{t}}(\cdot)$. Specifically, if $a_{t}^{*}\leq\tau$, then $p^{int^{*}}(a_{t}|\bar{y}_{t-1},a_{t}^{*},\bar{a}_{t-1},\bar{l}_{t},\bar{r}_{t},\bar{s}_{t}=1)=1,$ if $a_{t}=a_{t}^{*}$. However, if $a_{t}^{*}>\tau$, then
$p^{int^{*}}(a_{t}|\bar{y}_{t-1},a_{t}^{*},\bar{a}_{t-1},\bar{l}_{t},\bar{r}_{t},\bar{s}_{t}=1)=p^{\delta_{t}}(\delta_{t})I\{a_{t}=a_{t}^{*}-\delta_{t}\}.$
@@ 4 Intervention effects for mortal cohorts
Below we describe two types of mortal cohort inferences under the proposed incremental intervention on the NVT: partly conditional inference and inference on the principal strata. For simplicity, we temporarily restrict our description to the regimes $g\in(g^{int^{*}},g^{nc})$, where $g^{int^{*}}$ is the set of all deterministic intervention regimes based on the NVT, with support on the set of possible regimes $\mathcal{G}$, and $g^{nc}$ denote the natural course, i.e. the natural progression as if no intervention was considered. Furthermore, let $\bar{Y}_{T}(g)$, $\bar{A}_{T}(g)$, $\bar{R}_{T}(g)$, $\bar{S}_{T}(g)$, and $\bar{L}_{T}(g)$ represent the counterfactual outcome, treatment, missingness, survival, and covariate histories associated with the regime $g$.
@@ 4.1 Partly conditional inference
Partly conditional models have been proposed to address truncation by death under hypothetical interventions. For these models inference is conditioning on the sub-population being alive at a specific time-point, in here, being alive at end of follow-up, i.e. time $T$. Hence, inferences focus on what we call a partly conditional intervention effect ($PCIE$),
$PCIE=E[Y_{T}(g^{int^{*}})|S_{T}(g^{int^{*}})=1]-E[Y_{T}(g^{nc})|S_{T}(g^{nc})=1].$ (1)
Although the $PCIE$ gives a crude comparison of differences in potential outcomes under the two regimes, which is useful for prognostication, it often compares potential outcomes for different underlying populations. For example, it is likely that an sBP intervention may also affect survival rates. As such, the population surviving under the natural course will differ from the one surviving under the intervention.
@@ 4.2 Assumptions and identification of the PCIE
We now define a set of identifying assumptions for the $PCIE$.
@@
C2a Conditional exchangeability under the natural course: For $g^{nc}$ and $t=1,\ldots,T$, conditional on the history, the treatment and the potential outcome are independent. That is $\mkern 3.0mu\underline{\mkern-3.0muY\mkern-3.0mu}\mkern 3.0mu_{t}(g^{nc})\perp\!\!\!\perp A_{t}|\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$ and $\mkern 3.0mu\underline{\mkern-3.0muS\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{nc})\perp\!\!\!\perp A_{t}|\bar{Y}_{t}=\bar{y}_{t},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$.
@@
$\left(\mkern 3.0mu\underline{\mkern-3.0muY\mkern-3.0mu}\mkern 3.0mu_{t}(g^{int^{*}}),\mkern 3.0mu\underline{\mkern-3.0muA\mkern-3.0mu}\mkern 3.0mu_{t}^{*}(g^{int^{*}})\right)\perp\!\!\!\perp A_{t}|\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$ and $\left(\mkern 3.0mu\underline{\mkern-3.0muL\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{int^{*}}),\mkern 3.0mu\underline{\mkern-3.0muS\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{int^{*}})\right)\perp\!\!\!\perp A_{t}|\bar{Y}_{t}=\bar{y}_{t},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$. The assumption states that under the intervention, the NVT at time $t$ has no direct effect on the cognitive outcome, or survival, except through future treatment measures. Hence, an implication of this assumption is that the NVT is not itself a time-varying confounder [29, 40].
@@
C4 Missingness at follow-up: For the outcome, treatment and time-varying confounders we assume missing at random conditional on survival (MARS) on the distribution. In particular, for all $t>1$, $p(y_{t}|\bar{a}_{t},\bar{l}_{t},r_{t}=0,\bar{s}_{t}=1,\bar{y}_{t-1})=p(y_{t}|\bar{a}_{t},\bar{l}_{t},\bar{r}_{t}=1,\bar{s}_{t}=1,\bar{y}_{t-1})$, $p(a_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=0,\bar{s}_{t}=1,l_{t})=p(a_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=1,\bar{s}_{t}=1,l_{t})$, and $p(l_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=0,\bar{s}_{t}=1,l_{t-1})=p(l_{t}|\bar{y}_{t-1},\bar{a}_{t-1},r_{t}=1,\bar{s}_{t}=1,l_{t-1})$.
For the Betula data we perform a set of sensitivity analyses where we investigate robustness of results to violations of assumptions C2, C3, and C4. See Section 7.2.
@@
For longitudinal cohort data, under the assumptions C1, C2a, C3 and C4, the mean outcome at time $T$ under the natural course, $g^{nc}$, can be computed using the g-formula (Robins, 1986) conditioning on survival. That is
$\int_{\bar{s}_{T}}\int_{\bar{r}_{T}}\int_{\bar{l}_{T}}\int_{\bar{a}_{T}}\int_{\bar{y}_{T-1}}E[Y_{T}|\bar{y}_{T-1},\bar{a}_{T},\bar{l}_{T},\bar{r}_{T},\bar{s}_{T}=1]$
×$\quad\times\prod_{k=1}^{T}p(a_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{r}_{k},\bar{s}_{k}=1)$
×$\quad\times p(y_{k-1}|\bar{y}_{k-2},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1).$ (2)
To compute (S0.Ex2), Monte Carlo integration is implemented over the set of temporally preceding variables under the natural course, denoted by $H_{nc}$, among those who are alive at the final time point $\bar{s}_{T}=1$. Note, $s_{T}=1$ must also be sampled for.
@@
In contrast, for the intervention, $g^{int^{*}}$, under assumptions C1, C2b, C3 and C4, the mean outcome at time $T$ for $g^{int^{*}}$ can be computed using the Extended g-formula [31]. The main difference is that $p(a_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{r}_{k},\bar{s}_{k}=1)$ in (S0.Ex2) is replaced by $\int_{a_{t}^{*}}p^{int^{*}}(a_{k}|\bar{y}_{k-1},a^{*}_{k},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)\times p(a^{*}_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)$; details are given in Supplementary Appendix A. Similar to (S0.Ex2), Monte Carlo integration is implemented over the history under the intervention, denoted by $H_{int^{*}}$, among those who are alive. However, instead of sampling pseudo treatment data from the observed data distribution, as in (S0.Ex2), pseudo NVT data is first sampled from the observed data distribution. For those whose NVT does meet the threshold $\tau$, treatment is assigned to the NVT, and among those whose NVT does not meet the threshold $\tau$, treatment is assigned such that the NVT is shifted downwards by $\delta_{t}$, as described in Section 3.2.
@@
The $PCIE$ in (1) is then computed by subtracting the expected outcome under regime $g^{nc}$ from the expected outcome under regime $g^{int^{*}}$. More details on identification are found in Supplementary Appendix B.
4.3 Hypothetical interventions on the principal strata
Conditioning on survival as described in Section 4.1 may introduce bias due to the fact that survival is a post-randomization event [42]. Hence, the partly conditional approach described cannot properly address etiological research questions in mortal cohorts. One alternative estimand is the survivors average causal effect, i.e. the causal effect on the subpopulation surviving under both regimes, $g^{nc}$ or $g^{int^{*}}$, i.e. the principal strata ([8, 9]). Our goal here is to estimate the difference in expected potential outcomes between $g^{nc}$ and $g^{int^{*}}$ on the principal strata. That is, a survivors average intervention effect ($SAIE$),
$SAIE=E[Y(g^{int^{*}})-Y(g^{nc})\mid S_{T}(g^{nc})=S_{T}(g^{int^{*}})=1].$ (3)

### 20 Mar 2024 v4

@@
Summary In the present study we investigate overall population effects on episodic memory of an intervention over 15 years that reduces systolic blood pressure in individuals with hypertension. A limitation with previous research on the potential risk reduction of such interventions is that they do not properly account for the reduction of mortality rates. Hence, one can only speculate whether the effect is due to changes in memory or changes in mortality. Therefore, we extend previous research by providing both an etiological and a prognostic effect estimate. To do this, we propose a Bayesian semi-parametric estimation approach for an incremental threshold intervention, using the extended G-formula. Additionally, we introduce a novel sparsity-inducing Dirichlet hyperprior for longitudinal data, that exploits the longitudinal structure of the data. We demonstrate the usefulness of our approach in simulations, and compare its performance to other Bayesian decision tree ensemble approaches. In our analysis of the data from the Betula cohort, we found no significant prognostic or etiological effects across all ages. This suggests that systolic blood pressure interventions likely do not strongly affect memory, whether at the overall population level or in the population that would survive under both the natural course and the intervention (the always survivor stratum).
Key words: BART, dropout, G-computation, LDART, missing not at random, deaths
@@ 1. Introduction
Hypertension is a well-established risk factor for cognitive impairment and dementia [21]. However, the potential of promoting cognitive abilities through antihypertensive treatment remains poorly understood [11]. Studying the link between hypertension and cognition is complicated by the temporal dimension, where cognitive alterations caused by hypertension are processes that likely unfolds over decades rather than years, making it challenging to conduct randomized clinical trials (RCTs). Although prospective studies can never replace RCTs, applying more sophisticated statistical methods/methodology to existing datasets could provide additional insights.
@@
Two approaches for hypothetical interventions on a continuous risk factor, which only intervene on subjects that do not reach a prespecified threshold, are the threshold intervention [33] and the representative intervention (e.g., [24, 39]). The former assigns the threshold value to all individuals who do not meet the prespecified target, while the latter assigns a value drawn from the distribution among those in the population meeting the target level. Hence, both approaches ensure treatment remains within a prespecified interval. Nevertheless, it is common for a significant proportion of the population to be unable to control their sBP at optimal levels, due to factors such as poor response to medications. One alternative approach in these settings is an incremental intervention, where a subject’s treatment is instead shifted downward (or upwards) by some prespecified function [9, 22]. However, their approaches are not valid for longitudinal data with dropout and deaths.
Attrition due to dropout and deaths is common in longitudinal aging studies, especially when subjects are followed for an extended duration. The G-formula can address both missing response and missing covariate data under an assumption of missing at random (MAR; [28, 39, 16]). Further strategies for sensitivity analyses have been proposed in situations where the missingness is thought to be missing not at random (MNAR) [12, 13]. However, studying cognition after death is typically not defined and is not of interest, leading to consideration of a mortal cohort approach [4]. Mortal cohort inferences truncate subjects’ outcomes after death. The first approach, partly conditional inference (PCI), conditions on the sub-population that is still alive at the end of follow-up [17, 37, 36], while the second approach, principal stratification, conditions on the subpopulation that would survive through the end of follow-up regardless of treatment assignment, known as the always survivor stratum [6, 7]. Although the two approaches have similarities, they differ with respect to the type of information they generate. PCI has been proposed for estimating non-causal associations and is particularly useful for prognostication, i.e., when the interest is in the health outcome among all individuals who are alive. This approach has, for example, been considered through the use of the G-formula in hypothetical interventions targeting risk factors for coronary heart disease [33]. In contrast, principal stratification is useful in situations where survival differs between treated and control groups and has been applied when the interest is in etiological effect estimates (e.g. [10, 34, 35]).
@@
In the present study we investigate population effects on episodic memory of an sBP intervention using prospective cohort data from a large representative sample [23]. An important limitation in prior research is the lack of sufficient consideration of altered mortality rates when evaluating the potential risk reduction. As such, we propose both partly conditional inference and inference on the always survivor stratum. Thus, our study extends previous research by providing both etiological and prognostic intervention effect estimates, which are both critical for public health practice, offering a more comprehensive understanding of the intervention’s impact. For estimation, we propose a Bayesian semi-parametric approach based on the extended G-formula [27]. In addition, we introduce a novel sparsity-inducing Dirichlet hyperprior designed for regression with longitudinal data, and demonstrate the usefulness of our approach in a simulation relative to other Bayesian tree ensemble approaches.
@@
The paper is organized as follows. In Section 2, we introduce the Betula data. In Section 3, we introduce the proposed intervention in a Bayesian setting. In Section 4, we describe the two types of mortal cohort inferences. In Section 5, we provide computational details and introduce the novel sparsity-inducing Dirichlet hyperprior for longitudinal data. In Section 6, we conduct a simulation study to assess the operating characteristics of the new prior, and in Section 7, we apply the methodology to the Betula data to answer the substantive question of interest. We close in Section 8 with some conclusions.
@@ 2 A blood pressure intervention
The subjects in this study are participants in the Betula study, a population-based prospective cohort study initiated in 1988-1990, with the objective to study memory, aging and dementia [23]. We consider 15-years of longitudinal data collected over four time points, with data collection occurring at approximately five-year intervals. A total of n=800 participants were included in the current application, with 100 participants from each of the eight age cohorts $35,40,\ldots,70$ years old at enrollment. The cognitive outcome in this study was assessed using a composite episodic memory task, with a range of 0 to 76 ([14]). We included a set of baseline confounding factors, including age at enrollment, sex, and years of education, as well as a set of time-varying cardiovascular risk factors: body mass index (BMI), smoking, serum-cholesterol, and self-reported type 2 diabetes, as well as previous episodic memory measures. See Table A1 for baseline summary statistics. 38% of participants dropped out of the study before the 15-year follow-up, and among these, 18% died or developed dementia.
Although the recommended sBP levels have varied over time, we define hypertension as a sBP measure greater than 140 mm Hg. In Section 3 we propose an incremental threshold intervention. The intervention loosely imply that, at each follow-up time point, sBP levels is shifted downwards by a prespecified value in individuals with hypertension, while no intervention is imposed on individuals without hypertension.
3 Methods
Consider a prospective cohort study where measurements on a set of variables are collected on each of $i=1,2,\ldots,n$ subjects over a specified follow-up period at $t=1,2,\ldots,T$ fixed time points. Let $L_{it}$ denote the vector of time-varying confounders, and $A_{it}$ denote a continuous treatment, measured at time $t$. Denote the continuous outcome measure by $Y_{it}$, where interest in this study is the outcome measure at the last time point $T$. Furthermore, let $S_{it}=1$ indicate if a subject is alive at time point $t$; and $S_{it}=0$ otherwise. Let $R_{it}=1$ indicate if subject $i$ has completed the cognitive testing at time $t$; and $R_{it}=0$ otherwise. We have monotone missingness, so if $R_{it}=0$, then $R_{ik}=0$ for $k>t$. We let $T^{r}_{i}$ denote the number of time points a subject participates in the study, and $T^{s}_{i}\geq T^{r}_{i}$ denotes the number of time points the subject is alive. Let the observed history of a time-varying variable be denoted with an overbar and the future with an underbar. We also write $H_{i}$ to denote all the observed past history just prior to the outcome variable at the last time point $T$, such that $H_{i}=(\bar{Y}_{it-1}=\bar{y}_{it-1},\bar{A}_{it}=\bar{a}_{it},\bar{L}_{it}=\bar{l}_{it},\bar{R}_{it}=\bar{r}_{it},\bar{S}_{it}=1)$, with support on the set of possible histories $\mathcal{H}_{i}$ and where lower-case letters denote possible realizations of a random variable. At each time point $t$, we assume data is observed in the temporal ordering $S_{it},R_{it},L_{it},A_{it},Y_{it}$, and for each subject we observe $(\bar{s}_{iT^{s}_{i}},\bar{r}_{iT^{r}_{i}},\bar{l}_{iT^{r}_{i}},\bar{a}_{iT^{r}_{i}},\bar{y}_{iT^{r}_{i}})$. In what follows, we suppress the subscript $i$ to simplify notation.
3.1 A hypothetical incremental threshold intervention
@@ 4 Intervention effects for mortal cohorts
In this section we describe two types of mortal cohort inferences under the proposed incremental intervention on the NVT: partly conditional inference and inference on the the always survivor stratum. For simplicity, we temporarily restrict our description to the regimes $g\in\{g^{int^{*}},g^{nc}\}$, where $g^{int^{*}}$ is a deterministic intervention regime based on the NVT, with support on the set of possible regimes $\mathcal{G}$, and $g^{nc}$ denotes the natural course, i.e. the natural progression in the presence of no intervention. Let $\bar{Y}_{T}(g)$, $\bar{A}_{T}(g)$, $\bar{R}_{T}(g)$, $\bar{S}_{T}(g)$, and $\bar{L}_{T}(g)$ represent the counterfactual outcome, treatment, missingness, survival, and covariate histories associated with the regime $g$.
@@ 4.1 Partly conditional inference
Partly conditional models have been proposed to address truncation by death under hypothetical interventions. For these models inference is conditional on the sub-population alive at a specific time-point, in our application being alive at end of follow-up, i.e. time $T$. Hence, inferences focus on what we call a partly conditional intervention effect: $PCIE=E[Y_{T}(g^{int^{*}})|S_{T}(g^{int^{*}})=1]-E[Y_{T}(g^{nc})|S_{T}(g^{nc})=1]$. Although the $PCIE$ gives a crude comparison of differences in potential outcomes under the two regimes, which is useful for prognostication, it often implicitly compares potential outcomes for different underlying populations. For example, it is likely that an sBP intervention may affect survival rates; as such, the population surviving under the natural course will differ from the one surviving under the intervention.
@@ 4.2 Assumptions and identification of the PCIE
We now define a set of assumptions to identify the $PCIE$.
@@
C2a Conditional exchangeability under the natural course: For $g^{nc}$ and $t=1,\ldots,T$, conditional on the history, the treatment and the potential outcome are independent. That is $\mkern 3.0mu\underline{\mkern-3.0muY\mkern-3.0mu}\mkern 3.0mu_{t}(g^{nc})\perp\!\!\!\perp A_{t}|\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$ and
$\mkern 3.0mu\underline{\mkern-3.0muS\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{nc})\perp\!\!\!\perp A_{t}|\bar{Y}_{t}=\bar{y}_{t},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$.
@@
$\left(\mkern 3.0mu\underline{\mkern-3.0muY\mkern-3.0mu}\mkern 3.0mu_{t}(g^{int^{*}}),\mkern 3.0mu\underline{\mkern-3.0muA\mkern-3.0mu}\mkern 3.0mu_{t}^{*}(g^{int^{*}})\right)\perp\!\!\!\perp A_{t}|\bar{Y}_{t-1}=\bar{y}_{t-1},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$ and $\left(\mkern 3.0mu\underline{\mkern-3.0muL\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{int^{*}}),\mkern 3.0mu\underline{\mkern-3.0muS\mkern-3.0mu}\mkern 3.0mu_{t+1}(g^{int^{*}})\right)\perp\!\!\!\perp A_{t}|\bar{Y}_{t}=\bar{y}_{t},\bar{A}_{t-1}=\bar{a}_{t-1},\bar{R}_{t}=\bar{r}_{t},\bar{S}_{t}=1,\bar{L}_{t}=\bar{l}_{t}$.
This assumption states that under the intervention, the NVT at time $t$ has no direct effect on the cognitive outcome, or survival, except through future measurements of the treatment. An implication of this assumption is that the NVT is not itself a time-varying confounder [25, 38].
@@
C4 Missingness at follow-up: For the outcome, treatment and time-varying confounders, we assume missing at random conditional on survival (MARS). In particular, for all $t>1$, $p(y_{t}|\bar{a}_{t},\bar{l}_{t},r_{t}=0,\bar{s}_{t}=1,\bar{y}_{t-1})=p(y_{t}|\bar{a}_{t},\bar{l}_{t},\bar{r}_{t}=1,\bar{s}_{t}=1,\bar{y}_{t-1})$, and similarly for $a_{t}$ and $l_{t}$.
@@
For longitudinal cohort data, under the assumptions C1, C2a, C3 and C4, the mean outcome at time $T$ under the natural course, $g^{nc}$, can be computed using the g-formula (Robins, 1986) conditioning on survival,
$\int_{\bar{s}_{T}}\int_{\bar{r}_{T}}\int_{\bar{l}_{T}}\int_{\bar{a}_{T}}\int_{\bar{y}_{T-1}}E[Y_{T}|\bar{y}_{T-1},\bar{a}_{T},\bar{l}_{T},\bar{r}_{T},\bar{s}_{T}=1]\times\prod_{k=1}^{T}p(a_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{r}_{k},\bar{s}_{k}=1)$
×$\quad\times p(l_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k},\bar{s}_{k}=1)\times p(r_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k}=1)$
×$\quad\times p(s_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1)\times p(y_{k-1}|\bar{y}_{k-2},\bar{a}_{k-1},\bar{l}_{k-1},\bar{r}_{k-1},\bar{s}_{k-1}=1).$ (1)
For the Betula data we perform sensitivity analyses to investigate robustness of results to violations of assumptions C2, C3, and C4, (see Section 7).
@@
In contrast, for the intervention, $g^{int^{*}}$, under assumptions C1, C2b, C3 and C4, the mean outcome at time $T$ can be computed using the Extended g-formula [27]. The key difference is that $p(a_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{r}_{k},\bar{s}_{k}=1)$ in (S0.Ex1) is replaced by $\int_{a_{t}^{*}}p^{int^{*}}(a_{k}|\bar{y}_{k-1},a^{*}_{k},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)\times p(a^{*}_{k}|\bar{y}_{k-1},\bar{a}_{k-1},\bar{l}_{k},\bar{s}_{k}=1)$; details are given in Supplementary Appendix A. Similar to (S0.Ex1), Monte Carlo integration is implemented over the history under the intervention, denoted by $H_{int^{*}}$, among those who are alive. However, instead of sampling pseudo treatment data from the observed data distribution, as in (S0.Ex1), pseudo NVT data is first sampled from the observed data distribution. For those whose NVT does meet the threshold $\tau$, treatment is set to the NVT, and otherwise, treatment is assigned such that the NVT is shifted downwards by $\delta_{t}$, as described in Section 3.2.
@@
The $PCIE$ is then computed as the differences in expected outcomes under the two regimes $g^{nc}$ and $g^{int^{*}}$. More details are given in Supplementary Appendix B.
4.3 Inference on the always survivor stratum